April 25-27

– Three stores provided ~39.25 hours of luminosity
– Pbar personnel repaired an LCW leak in Transport
– Linac personnel brought the I Source online
– Linac RF tank #2 lost its vacuum due to a bad pickup ceramic
– Tevatron personnel conducted PIXEL detector studies
– Pbar personnel traced emittance blowups to dampers
– A1 beam transfer line suffered vacuum burst
– Operators filled MI kickers with Fluorinert

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts